OpenStudy (turingtest):

I am getting a headache trying to figure out if wolf's answer to this integral is even correct.

OpenStudy (turingtest):

\[\int\csc^{10}\cot^3xdx\]I can do this integral in about 4 steps, but wolf seems to think you need about 100 http://www.wolframalpha.com/input/?i=integral%28csc%28x%29%29%5E%2810%29%28cot%28x%29%29%5E3dx the answers look similar, but the trig identities wolf applied mess me up can anyone verify wolf's answer?

OpenStudy (turingtest):

nevermind, working it backwards wolf gives the same answer as I get what a ridiculous way to do the problem though !

OpenStudy (turingtest):

I was hoping wolf was wrong, but this time it was just stupid

OpenStudy (anonymous):

Wolfram isn't probably wrong, but I am sure you're neither. Sometimes wolfram is just useless :/

OpenStudy (turingtest):

Yes, please click "show steps" to see why I had my doubts about Wolfram's answer whenever it goes on too long it tends to make more mistakes

OpenStudy (anonymous):

lol *scrolling*

OpenStudy (amistre64):

since the wolf uses algorithms to attempt to apply rules to a problem, its not going to necessarily go the human route. It will pick one option and work it out the best its been designed to do.

OpenStudy (turingtest):

Here's my way\[\int\csc^{10}x\cot^3xdx=\int\csc^9x\cot^2x(\csc x\cot x)dx\]\[=\int\csc^9(\csc^2-1)(\csc x\cot x)dx=\int(\csc^{11}x-\csc^9x)(\csc x\cot x)dx\]\[=\frac1{10}\csc^{10}x-\frac1{12}\csc^{12}x+C\]the heck with simplifying